The minimum distance required to stop a car traveling at 35 mi/hr is 40 ft. Assuming the same acceleration, what is the minimum distance required to stop a car traveling at 75 mi/hr?
1
Expert's answer
2019-05-16T02:44:39-0400

Let's first find the deceleration of the car from the kinematic equation:


"v_f^2 = v_i^2 + 2as,"

here, "v_f = 0" is the final velocity of the car, "v_i = 35 mi\/hr" is the initial velocity of the car, "a" is the deceleration of the car and "s" is the minimum distance traveled by the car until it stops.

Then, from this formula we can find the deceleration of the car:


"a = -\dfrac{v_i^2}{2s} = -\dfrac{(35 \dfrac{mi}{hr} \cdot \dfrac{5280 ft}{1 mi})^2}{2 \cdot 40 ft} = -4.27 \cdot 10^8 \dfrac{ft}{hr^2}."

The sign minus indicates that the car decelerates.

Finally, we can find the minimum distance required to stop a car traveling at 75 mi/hr assuming the same rate of deceleration:


"v_f^2 = v_i^2 + 2as,""s = -\dfrac{v_i^2}{2a} = -\dfrac{(75 \dfrac{mi}{hr} \cdot \dfrac{5280 ft}{1 mi})^2}{2 \cdot (-4.27 \cdot 10^8 \dfrac{ft} {hr^2})} = 184 ft."

Answer:

"s = 184 ft."